write an equation of the line that passes through each point with the given slope.
(-3,-5),m=-5/3(5 over 3)

Answer by venugopalramana(3286)

(-3,-5),m=-5/3(5 over 3)
EQUATION OF A LINE PASSING THROUGH POINT (X1,Y1)WITH SLOPE M IS GIVEN BY
Y-Y1=M(X-X1)...SO...
Y-(-5)=(-5/3){X-(-3)}
Y+5=(-5/3)(X+3)........MULTIPLYING WITH 3 THROUGHOUT
3Y+15=-5X-15
5X+3Y+30=0